OpenStudy (anonymous):

For which equation is x = 4 a solution? A. 5(3 – 2x) = 7 B. –3(x – 2) = –6 C. 2(x – 5) = 3 D. 4(3x + 1) = 49

OpenStudy (anonymous):

B. -3(x-2) = -6 -3x + 6 = -6 -6 -6 -3x = -12 -------- -3 -3 x = 4 negative and negative when multiplied equals a positive